OpenStudy (anonymous):

pls answer this multiplication of polynomials. the sign of that * is in the nxt of exponent.. (10xy*5)(3x*3y)(-3X*4z*3) whats the answer??/ plss

OpenStudy (anonymous):

\[(10x y*5)(3x*3y)(-3x*4z*3)=-16200 x^3 y^2 z \]

OpenStudy (anonymous):

how can u get that?

OpenStudy (anonymous):

Gather like terms. You can replace the inner parentheses combination ")(" with a "*". (10*4*3*3*(-3)*4*3)*(x*x*x)*(y*y)*(z) Starting from left to right: (10*4*3*3*(-3)*4*3) = -16200 (x*x*x) = x^3 (y*y) = y^2 z = z Does that help you?

jhonyy9 (jhonyy9):

if this sign * is the exponent sign you will get (-90x*(8)y*(6)z*(8) ) --- because when you multiply x*2 with x*3 the exponents you need to assumes and hence you will get x*5 - so hope that you have understood it

OpenStudy (anonymous):

The ^ sign is the conventential exponent sign. * is the multiply sign and has the same meaning as the x sign.